# Aeos: Star-by-Star Cosmological Simulations of Early Chemical Enrichment and Galaxy Formation

### v1

@@ 4.3 Metallicity Floors from Pop III Enrichment
Our simulation supports a Pop III metallicity floor of Z = 10-4 Z that generally holds for halo masses of $10^{5}$ to $10^{7}$ M. At the highest halo masses, the metallicity of the gas increases, but there are only 4 halos above 5×$\times 10^{6}$ M, so this may or may not be robust. The best-fit metallicity floor of each element is approximately: [C/H] = -4.1, [N/H] = -5.3, [O/H] = -4.0, [Na/H] = -4.9, [Mg/H] = -4.4, [Ca/H] = -4.9, [Mn/H] = -5.4, and [Fe/H] = -5.1. There is significant scatter, however, due to inhomogeneous mixing and explosive outflows. We exclude Sr and Ba as they are not produced in significant amounts by Pop III.
We also look at the stellar chemical abundances of the first low-mass Pop II stars in each galaxy to see how the gas enrichment translates into stellar enrichment for the first generation Pop II stars. 17 galaxies have begun Pop II star formation (see Figure 3), and we plot the chemical abundances of the first-generation low-mass Pop II stars from each Pop II galaxy in Figure 9. These abundances are similar to the gas metallicity floors in Figure 8, as we would expect, with some differences (e.g., some of the median [X/H] abundances are noticeably lower) that are likely due to the smaller sample size and inohomogeneous metal mixing. As seen by the error bars, the scatter in the median chemical abundances is over a dex, showing significant variation in the abundances of the first generation of Pop II stars across different galaxies. The median [O/H] of the first generation of low-mass Pop II stars is [O/H] = -4.1, consistent with that of the Pop III halo gas.
@@ 5 Conclusions
Additionally, we estimate metallicity floors for different metals (e.g., [O/H] = -4) as a result of Pop III enrichment. These metallicity floors are fairly flat even with variations in halo mass and time, though with significant scatter. There is also over a dex of scatter in the median metallicity of first generation Pop II stars from different galaxies, showing that significant variation exists between first generation Pop II. We also investigate CEMP signatures in the first generation of Pop II stars and compare to the frequencies of CEMP stars observed in the Milky Way.
@@ Appendix A Metal Yields
Figure A.1 illustrates the mass yields used in the Aeos simulations for ten tracked elements—C, N, O, Na, Mg, Ca, Mn, Fe, Sr, and Ba—across different stellar progenitor masses, highlighting contributions from various nucleosynthetic sources. These sources include Pop III core-collapse supernovae (CCSN), Pop II CCSN, AGB winds, SNIa, and massive star winds. Each curve represents the mass yield of an element as a function of progenitor mass, revealing the variations in yield patterns due to different evolutionary processes and initial conditions. The plot emphasizes the mass-dependent nature of element production, crucial for understanding the chemical evolution in our simulations. For more information, see Section 2.4.

### v2

@@ 4.3 Metallicity Floors from Pop III Enrichment
Our simulation supports a Pop III metallicity floor of Z = 10-4 Z that generally holds for halo masses of $10^{5}$ to $10^{7}$ M. At the highest halo masses, the metallicity of the gas increases, but there are only 4 halos above 5×$\times 10^{6}$ M, so this may or may not be robust. The best-fit metallicity floor of each element is approximately: [C/H] = -4.1, [N/H] = -5.3, [O/H] = -4.0, [Na/H] = -4.9, [Mg/H] = -4.4, [Ca/H] = -4.9, [Mn/H] = -5.4, and [Fe/H] = -5.1. There is significant scatter, however, due to inhomogeneous mixing that follows localized enrichment from different SNe and explosive outflows. We exclude Sr and Ba as they are not produced in significant amounts by Pop III.
We also look at the stellar chemical abundances of the first low-mass Pop II stars in each galaxy to see how the gas enrichment translates into stellar enrichment for the first generation Pop II stars. 17 galaxies have begun Pop II star formation (see Figure 3), and we plot the chemical abundances of the first-generation low-mass Pop II stars from each Pop II galaxy in Figure 9. These abundances are similar to the gas metallicity floors in Figure 8, as we would expect, with some differences (e.g., some of the median [X/H] abundances are noticeably lower) that are likely due to the smaller sample size and inhomogeneous metal mixing. As seen by the error bars, the scatter in the median chemical abundances is over a dex, showing significant variation in the abundances of the first generation of Pop II stars across different galaxies. The median [O/H] of the first generation of low-mass Pop II stars is [O/H] = -4.1, consistent with that of the Pop III halo gas.
@@ 5 Conclusions
Additionally, we estimate metallicity floors for different metals (e.g., [O/H] = -4) as a result of Pop III enrichment. These metallicity floors are fairly flat even with variations in halo mass and time, though with significant scatter. There is also over a dex of scatter in the median metallicity of first generation Pop II stars from different galaxies, showing that significant variation exists between first generation Pop II. We also investigate CEMP signatures in the first generation of Pop II stars and compare to the frequencies of CEMP stars observed in the Milky Way. [C/Fe] drops with metallicity in the simulations as it does in the observations.
@@ Appendix A Metal Yields
Figure A.2: The ratio of the yields of each metal from each astrophysical source in the simulation, with respect to iron (Fe). The shaded regions include the yield range for progenitor stars at different metallicities for a given progenitor mass.
Figure A.1 illustrates the mass yields used in the Aeos simulations for ten tracked elements — C, N, O, Na, Mg, Ca, Mn, Fe, Sr, and Ba — across different stellar progenitor masses, highlighting contributions from various nucleosynthetic sources. These sources include Pop III core-collapse supernovae (CCSN), Pop II CCSN, AGB winds, SNIa, and massive star winds. Each curve represents the mass yield of an element as a function of progenitor mass, revealing the variations in yield patterns due to different evolutionary processes and initial conditions. The plot emphasizes the mass-dependent nature of element production, crucial for understanding the chemical evolution in our simulations. For more information, see Section 2.4. We also show the ratio of the yields of each metal species with respect to Fe (Figure A.2).
The yield sets used for this study were chosen based on a combination of factors, including alignment with observed data and ease of implementation in simulations. For AGB stars (M < 8 M), we selected the yields from Cristallo et al. (2015), which were recommended for their ability to capture the production of elements like Sr and Ba. For massive stars (M > 8 M), we adopted the yields from Limongi & Chieffi (2018) (LC18), as they offer a detailed treatment of stellar rotation and its impact on yields, despite some challenges like the underproduction of Mg. The LC18 yields also assume that stars above 25 M undergo direct collapse without producing supernova yields, a feature consistent with previous assumptions in other models used by the team. This choice balances the need for physical realism with practical considerations of implementing yield tables that separate contributions from stellar winds and supernovae, allowing for a more flexible simulation setup. Despite the challenges of managing memory usage with many tracer fields, this combination of yield sets was found to produce the most realistic results when compared with Milky Way observations in galactic chemical evolution (GCE) models using OMEGA (Côté & Ritter, 2018), making it the preferred choice for the study.
The decision to increase the magnesium (Mg) yields by a factor of 2.2 was made to resolve a mismatch between the theoretical yields predicted by the LC18 model and the observed [Mg/Fe] ratios in metal-poor stars of the Milky Way. The unmodified LC18 yields tended to underproduce Mg, leading to lower-than-expected [Mg/Fe] ratios in simulations that track the chemical evolution of galaxies. Observational data from stars with [Fe/H] < -2 indicate that Mg should be more abundant, reflecting the contributions of core-collapse supernovae in the early Universe. The 2.2x scaling factor was derived through iterative comparisons between simulated results using our chosen yields with OMEGA (Côté & Ritter, 2018) and stellar chemical abundances of metal-poor stars in the Milky Way from JINAbase (Abohalima & Frebel, 2018), effectively bringing the simulated [Mg/Fe] ratios into alignment with what is observed (see Figure A.3). This increase in Mg does not affect gas cooling in our simulation because it does not provide significant cooling (e.g., Smith et al., 2017) and in Aeos the gas cooling is determined by total metallicity. While the Mg adjustment is empirical and does not address the physical processes behind the underproduction, it ensures that the simulations more accurately reflect the chemical enrichment history of early stellar populations, improving the overall consistency of the model with observed data.
Figure A.3: Top: Shown as lines, simulated stellar chemical abundances for different stellar rotation velocities in a galactic chemical evolution OMEGA model using AGB wind yields from Cristallo et al. (2015) and Pop II CCSN and massive star wind yields from Limongi & Chieffi (2018). In orange, observations of metal-poor stars in the Milky Way. Bottom: Results after boosting the Mg yields from LC18 by a factor of 2.2 to match observations.